. A motor with coil of area 3cm^2 rotes with uniform angular velocity w in the magnetic field changing at 200 T per minute. A) make a graphical representation of induced EMF; b) find such w.


1
Expert's answer
2021-12-06T17:45:03-0500

"\varepsilon=\frac{ \Phi }T=S\frac BT=1~mV,"

"\omega=\frac{2\pi}T=0.105~\frac{rad}s."